# Local Topological Constraints on Berry Curvature in Spin–Orbit Coupled BECs

### v1

@@ 9 Conclusions
We have shown that a two-component spin–orbit-coupled BEC possesses a built-in geometric obstruction preventing the simultaneous flattening of Berry curvature along the two phase directions $\phi_{+}$ and $\phi_{-}$, even when all global Chern numbers vanish. The extended parameter space
$M=T^{2}_{\mathrm{BZ}}\times S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}},$
equipped with the natural Kaluza–Klein metric induced by the synthetic Berry connections, carries a metric connection $\nabla^{C}$ whose torsion encodes the synthetic gauge fields of the condensate. The harmonic part of this torsion defines a mixed cohomology class
$\omega]\in H^{2}(T^{2}_{\mathrm{BZ}})\otimes\bigl(H^{1}(S^{1}_{\phi_{+}})\oplus H^{1}(S^{1}_{\phi_{-}})\bigr),$
with mixed tensor rank $r=1$ for any generic SOC configuration in which the two Berry curvature classes are both non-zero. Because $\dim H^{2}(T^{2}_{\mathrm{BZ}})=1$, the two Berry curvature classes are necessarily cohomologically proportional, forcing the mixed class to be a single simple tensor. Moreover, the Kaluza–Klein geometry admits no parallel one-forms along the phase directions, so the obstruction kernel vanishes, $\mathcal{K}=0$, giving the reduced rank $r^{\sharp}=1$.
Applying the PT lower bound for connections with skew torsion yields the sharp inequality
@@
This algebraic constraint on the curvature implies that the momentum degrees of freedom cannot be decoupled from the two-phase directions by any smooth gauge transformation. Importantly, the obstruction measured by $r^{\sharp}$ does not depend on the total Chern number: it persists even in regimes where global Berry flux cancels, a situation in which conventional topological invariants predict no restriction. The Rashba–Dresselhaus example illustrates compatibility with familiar Chern-number phenomena (e.g., $|c_{1}|=1$ in the simplest case), while also highlighting that the mixed-rank bound continues to operate beyond the Chern-number paradigm.
Experimentally, the bound suggests interferometric protocols capable of resolving the independent Berry phases associated with $\phi_{+}$ and $\phi_{-}$, thereby detecting the locally irremovable curvature. Conceptually, it identifies a new mechanism for topological robustness in synthetic gauge systems—one controlled not by global integrals of curvature but by the algebraic structure of its mixed components. We expect this framework to extend naturally to multicomponent cold-atom systems and synthetic-dimensional lattices where multiple phase or internal degrees of freedom couple to momentum-space geometry.

### v2

@@ 9 Conclusions
We have demonstrated that a two-component spin–orbit-coupled (SOC) BEC possesses an intrinsic geometric obstruction that prevents the simultaneous flattening of Berry curvature along the independent phase directions $\phi_{\pm}$. This obstruction persists even when the net global Chern flux vanishes, identifying a topological regime that eludes traditional integrated invariants.
The extended parameter space $M=T^{2}_{\mathrm{BZ}}\times S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}$, equipped with the natural Kaluza–Klein metric $g_{M}$ induced by synthetic gauge fields, carries a metric connection $\nabla^{C}$ with totally skew-symmetric torsion. The harmonic part of this torsion represents a mixed cohomology class $\omega]$ with mixed tensor rank $r=1$ for any generic configuration where the individual Berry curvature classes are non-trivial. Due to the one-dimensionality of $H^{2}(T^{2}_{\mathrm{BZ}})$, these classes are necessarily proportional, forcing $\omega]$ to take the form of a single simple tensor. Crucially, we have shown that the non-vanishing Berry curvature acts as a geometric obstruction to the integrability of the horizontal distribution, ensuring the vanishing of the obstruction kernel $\mathcal{K}=0$. Consequently, the bound $r^{\sharp}=1$ holds directly for the physical metric $g_{M}$.
Applying the PT lower bound yields the sharp inequality:
@@
This algebraic constraint implies that momentum degrees of freedom are fundamentally “locked” to the phase sectors, an irreducible coupling that cannot be removed by smooth gauge transformations. The persistence of $r^{\sharp}$ in regimes with zero total Chern number—as illustrated by our refined Rashba–Dresselhaus analysis—highlights that the mixed-rank bound operates beyond the Chern-number paradigm, detecting structural features of the gauge field that integrated invariants fail to resolve.
Experimentally, this framework suggests interferometric protocols to resolve independent Berry phases, providing a direct signature of locally irremovable curvature. Conceptually, it establishes a new mechanism for topological robustness in synthetic gauge systems, controlled by the algebraic structure of mixed cohomology components. We expect this framework to be readily applicable to multicomponent cold-atom systems and synthetic-dimensional lattices, where multiple internal degrees of freedom couple to momentum-space geometry.